OpenStudy (anonymous):

What is the derivative of 1/x? Is it -2?

OpenStudy (mimi_x3):

-1/x^2 Where did you get -2 from xD

OpenStudy (anonymous):

1/x =\[x ^{-1}\] = \[x ^{-1-1}\] = \[x ^{-2}\]? lol

OpenStudy (mimi_x3):

http://www.wolframalpha.com/input/?i=1%2Fx

OpenStudy (mimi_x3):

And the way your doing it looks wrong already (:

OpenStudy (anonymous):

I kind of get it, but where did they negative sign come from? lol

OpenStudy (mimi_x3):

Well , 1/x =x^-1

OpenStudy (anonymous):

*the

OpenStudy (mimi_x3):

Do you know how to do it now ?

OpenStudy (anonymous):

Yeah, It's confusing using the short cut so I found it with the difference quotient, and the answer is the same.

OpenStudy (anonymous):

1/x =x^-1 ===> d/dx=-x^-2 = -1/x^2